2 Let x equal a real number that is selected randomly from the closed interval from zero to ten. Use your intuition to assign values to

  1. P{x: 0≤x≤3}=____________

  1. P{x: x=3}=________________
  1. P{x: 0≤x≤10}=_____________

  1. P{x: x≥7}=_______________
Answer #1

for this follows uniform distribution

a) P{x: 0≤x≤3}= (3-0)/10=0.3

b) P{x: 0≤x≤10}= (10-0)/10 =1

c) P{x: x≥7}= (10-7)/10=0.3